INAF queue observing, 2015 Apr 26-27

Observer: F. Cusano, R. Carini
Telescope Operator: Geno Bechetti
Support Astronomer: Berry Rothberg
Instrument: LUCI

Summary

There is snow, the dome is closed.

Night duration (12\xB0 el twilight) 20:00-04:30 (3:00-11.30 UT) = 8 hrs 30 m

Observing time: 0 hrs

Weather time loss: 8.5 hrs (100%)

Data Summary

There is snow, the dome is closed

Project_RUN Objects Filter ExpTime Notes

LUCI Calibration

Details

04:50 we start LUCI calibrations
  • luci.20150427.0001-0005 dark tests
  • start with Piconcelli Program, executeLUCIScript.sh A5_LS_Ks.flat luci.20150427.0006-0015
  • executeLUCIScript.sh A5_LS_Ks.arc luci.20150427.0016-0019
  • A5 and A6 have the same configuration
  • executeLUCIScript.sh B1_LS_210_H.flat luci.20150427.0020-0029
  • executeLUCIScript.sh B1_LS_210_H.arc luci.20150427.0030-0033
  • B1 and B2 have the same configuration
  • Della Caca Program, executeLUCIScript.sh Target_11D.flat luci.20150427.0034-0043
  • executeLUCIScript.sh Target_11D.arc luci.20150427.0044-0047
  • D, F , G have the same configuration
  • Rossi program
  • executeLUCIScript.sh GRB070521_ls_200_H+K_zJ.flat luci.20150427.0048-0067
  • executeLUCIScript.sh GRB070521_ls_200_H+K_zJ.arc luci.20150427.0068-0079
  • executeLUCIScript.sh GRB070521_ls_200_H+K_HK.flat luci.20150427.0080-0099
  • executeLUCIScript.sh GRB070521_ls_200_H+K_HK.arc luci.20150427.0100-0107
  • back to Della Ceca program
  • executeLUCIScript.sh Target_11E.flat luci.20150427.00108-00117
  • executeLUCIScript.sh Target_11E.arc luci.20150427.00118-00123

-- %USERSIG{RobertaCarini - 2015-04-27}%

Comments

 
Topic revision: r7 - 29 Apr 2015, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback